# Auto-Search and Refinement: An Automated Framework for Gender Bias Mitigation in Large Language Models

### v1

@@ Appendix A Auto-search Algorithm
This process is repeated for every position in the Fairwords, resulting in $b\times l$ potential substitutions. From these, we randomly sample $k$ Fairwords as candidates, denoted as $\mathcal{K}=\{s^{\prime}\}$, and compute the exact loss for each candidate using Equation A. The token replacement that minimizes the loss is chosen as the final replacement.
The entire auto-search procedure is outlined in the pseudo-code provided in Algorithm 1. In our experiment, we set the length of the Fairwords $l$ to 20, the batch size $m$ to 25, the number of top-weight candidates $b$ to 256, and the sampling size $k$ to 512. All searches are performed on a single NVIDIA A40 GPU, with the optimization process taking around 24 hours to complete 300 steps.
@@
$\mathtt{current\_Fairwords}=[\mathtt{random\_init\_a\_Fairwords()}]$;
@@
[…]
       $[(x^{(j)},y_{c}^{(j)}),y_{r}^{(j)})]_{j=1\ldots m}\sim\mathcal{D}$;
@@
[…]
@@
[…]
@@
[…]
@@
Appendix B Dataset Examples
We use both gender-relevant and general tasks for a comprehensive assessment. The GA-test and BBQ-gender are used for gender-relevant tasks, while Dolly Eval, Instruct Eval, and BPO Eval are used for general tasks. Detailed descriptions of these datasets are provided below:
@@ Appendix B Dataset Examples
  • BBQ-gender consists of gender identity queries from the standard multiple-choice bias benchmark BBQ Parrish et al. (2021). Each BBQ question has an ambiguous and disambiguated version, with the latter adding context to enable the correct answer, allowing BBQ to assess LLMs’ gender bias in different contexts.
  • Dolly Eval samples 200 human-generated instructions across various task categories from dolly dataset Conover et al. (2023).
  • Instruct Eval Wang et al. (2022) consists of 252 expert-written tasks and instructions designed to assess the instruction-following capabilities of LLMs in user-oriented applications.
  • BPO Eval, created by Cheng et al. (2023), consists of 200 queries sampled from four open-source prompt datasets: OASST1, HH-RLHF, Chatbot Arena, and Alpaca-GPT4.
@@
Appendix C Additional Experiment results
C.1 Win-tie-loss on GA-test
Figure 7 shows the performance comparison between the base models and the models after applying our method, with Llama3.1 as the evaluator. The same trend is observed, where responses generated after applying FaIRMaker consistently outperform the original ones.
Figure 7: Performance comparison between the base models and the models after applying FaIRMaker on the GA-test dataset, with Llama3.1 as the evaluator.
C.2 RS on GA-test
Table 8 shows the average RS for each LLM, with the highest scores highlighted in bold, as evaluated by Llama3.1. FaIRMaker consistently outperforms other baseline methods across both white-box and API-access models, demonstrating its strong capability in dialogue generation.
Model Llama3.1 Score ($\uparrow$)
Ori. FM. Interv. CF-D Desc-D
Llama2-Alpaca 4.38 4.68 4.58 4.16 ($\downarrow$) 4.06 ($\downarrow$)
Llama2-Chat 4.92 4.94 4.89 4.12 ($\downarrow$) 4.53 ($\downarrow$)
Qwen2-Instruct 4.92 4.96 4.96 4.71 ($\downarrow$) 4.85 ($\downarrow$)
Qwen2.5-Instruct 4.93 4.98 4.96 4.47 ($\downarrow$) 4.50 ($\downarrow$)
GPT3.5-turbo 4.94 4.97 4.97 4.73 ($\downarrow$) 4.94
@@ C.2 RS on GA-test
C.3 Extending FaIRMaker with DPO
Figure 8 shows the overall performance of the model enhanced with DPO, FaIRMaker, and their combination, evaluated by Llama3.1. Although the scores vary, both evaluators exhibit the same trends that demonstrate the extendability of FaIRMaker.
Figure 8: Overall performance of FaIRMaker, DPO and their combination, evaluated by Llama3.1
C.4 Results of FaIRMaker w/o refinement
@@ C.5 Interpretability
Appendix D Prompts
D.1 Baseline Prompt
@@
Intervention. We should treat people from different socioeconomic statuses, sexual orientations, religions, races, physical appearances, nationalities, gender identities, disabilities, and ages equally. When we do not have sufficient information, we should choose the unknown option, rather than making assumptions based on our stereotypes.
CF-D. Despite being a male, Austin became a dental hygienist.
Desc-D. Dental hygienists focus on promoting oral health and hygiene.
D.2 Evaluation Prompt
@@
D.3 Refined Fairwords Construction Prompt

### v2

@@ Appendix A Auto-search Algorithm
This process is repeated for every position in the Fairwords, resulting in $b\times l$ potential substitutions. From these, we randomly sample $k$ Fairwords as candidates, denoted as $\mathcal{K}=\{s^{\prime}\}$, and compute the exact loss for each candidate using Equation A. The token replacement that minimizes the loss is chosen as the final replacement. The entire auto-search procedure is outlined in the pseudo-code provided in Algorithm 1.
@@
[…]
@@
  $\mathtt{candidate\_list}=\text{empty list}$;
  $\mathtt{Fairwords\_list}=\text{empty list}$;
  $(x^{(j)},y_{c}^{(j)}),y_{r}^{(j)})]_{j=1\ldots m}\sim\mathcal{D}$;
@@
     $\mathtt{candidate\_words}=\text{get }b\text{ words with maximum }\mathtt{weight}$;
@@
    $\mathtt{new\_candidates}=\text{random choose }k\text{ in }\mathtt{candidate\_list}$;
@@
 $\mathtt{best}\leftarrow\mathtt{Fairwords\_list}\text{ with minimize loss}$ ;
@@
Appendix B Detailed Experiment Configuration
Fairwords searching setup. In our experiment, we set the length of the Fairwords $l$ to 20, the batch size $m$ to 25, the number of top-weight candidates $b$ to 256, and the sampling size $k$ to 512. All searches are performed on a single NVIDIA A40 GPU, with the optimization process taking around 24 hours to complete 300 steps.
Refiner training setup. We fine-tune the Llama3.2-3B-Instruct model using LoRA with rank $r=8$, scaling factor $\alpha=16$, and a dropout rate of 0.05. The optimization is performed using the AdamW optimizer with a learning rate of $1\mathrm{e}{-5}$ and a cosine learning rate scheduler. Training is conducted for 10 epochs with 4-step gradient accumulation to simulate a larger batch size, and all experiments are run in FP16 precision for efficiency.
Appendix C Dataset Description
In this work, we utilize publicly available datasets for training and evaluating FaIRMaker, including GenderAlign [49], Self-Instruct [42], and BPO Eval [7] (under Apache License), as well as BBQ [30], Alpaca [37], and Free Dolly [9] (under Creative Commons License). Some of these datasets contain content that may be offensive or distressing. We assert that these data are solely used for the purpose of mitigating gender bias and improving model performance. We use both gender-relevant and general tasks for a comprehensive assessment. The GA-test and BBQ-gender are used for gender-relevant tasks, while Dolly Eval, Instruct Eval, and BPO Eval are used for general tasks. Detailed descriptions of these datasets are provided below:
@@ Appendix C Dataset Description
  • BBQ-gender consists of gender identity queries from the standard multiple-choice bias benchmark BBQ [30]. Each BBQ question has an ambiguous and a disambiguated version, with the latter providing additional context such that one answer becomes correct, which is typically the one against the stereotype. For gender-related questions, the answer choices include a male subject, a female subject, and unknown. sDIS is the bias score in disambiguated contexts, where one answer is correct and bias can be measured based on the model’s preference. It can be defined as: $\text{sDIS}=\#biased\_answer/\#non-unknown\_answer$. sDIS measures the proportion of biased answers among all non-unknown responses. A score of 0 indicates no gender bias. sAMB is the bias score in ambiguous contexts, where the correct answer is always “unknown”, and any gendered answer may reflect biased reasoning. It is computed as: $\text{sAMB}=(1-Accuracy)\times\text{sDIS}$. This formulation penalizes biased predictions more when the model is confidently incorrect. Lower sAMB scores indicate fairer behavior in uncertain scenarios.
  • Dolly Eval samples 200 human-generated instructions across various task categories from dolly dataset [9].
  • Instruct Eval [42] consists of 252 expert-written tasks and instructions designed to assess the instruction-following capabilities of LLMs in user-oriented applications.
  • BPO Eval, created by Cheng et al. [7], consists of 200 queries sampled from four open-source prompt datasets: OASST1, HH-RLHF, Chatbot Arena, and Alpaca-GPT4.
@@
Appendix D Additional Experiment results
D.1 RS on GA-test
Table 9 shows the average RS for each LLM, with the highest scores highlighted in bold, as evaluated by GPT4 and Gemini. FaIRMaker consistently outperforms other baseline methods across both white-box and API-access models, demonstrating its strong capability in dialogue generation.
Model GPT4 Score ($\uparrow$) Gemini Score ($\uparrow$)
Ori. FM. Interv. CF-D Desc-D MBIAS Ori. FM. Interv. CF-D Desc-D MBIAS
Llama2-Alpaca 3.27 3.77 3.68 3.09 ($\downarrow$) 2.94 ($\downarrow$) 3.21 ($\downarrow$) 3.48 3.99 3.97 3.22 ($\downarrow$) 3.03 ($\downarrow$) 3.23 ($\downarrow$)
Llama2-Chat 4.47 4.73 4.47 3.89 ($\downarrow$) 3.89($\downarrow$) 4.27 ($\downarrow$) 4.62 4.72 4.58 ($\downarrow$) 3.74 ($\downarrow$) 3.94 ($\downarrow$) 4.04 ($\downarrow$)
Qwen2-Instruct 4.58 4.81 4.74 4.34 ($\downarrow$) 4.34($\downarrow$) 4.36($\downarrow$) 4.77 4.84 4.84 4.54 ($\downarrow$) 4.50 ($\downarrow$) 4.56 ($\downarrow$)
Qwen2.5-Instruct 4.68 4.88 4.82 4.21 ($\downarrow$) 4.00 ($\downarrow$) 4.60 ($\downarrow$) 4.79 4.88 4.87 4.38 ($\downarrow$) 4.18 ($\downarrow$) 4.73 ($\downarrow$)
GPT3.5-turbo 4.72 4.88 4.87 4.60 ($\downarrow$) 4.60 ($\downarrow$) 4.59 ($\downarrow$) 4.92 4.96 4.81 ($\downarrow$) 4.80 ($\downarrow$) 4.96 4.89($\downarrow$)
@@ D.1 RS on GA-test
D.2 RS on Instruction Following Tasks
Table 10 shows the average RS across three instruction following datasets for each LLM, with the highest scores highlighted in bold, as evaluated by GPT4 and Llama3.1. FaIRMaker consistently outperforms other baseline methods across both white-box and API-access models, demonstrating its strong capability in instruction following tasks.
Model GPT4 Score
Dolly Eval Instruct Eval BPO Eval
Ori. FM. Ori. FM. Ori. FM.
Llama2-Alpaca 1.96 2.96 3.88 4.06 2.25 2.81
Llama2-Chat 3.92 3.93 4.01 4.08 3.71 4.40
Qwen2-Instruct 4.55 4.57 4.58 4.59 4.53 4.54
Qwen2.5-Instruct 4.52 4.47 ($\downarrow$) 4.80 4.78 ($\downarrow$) 4.51 4.51
GPT3.5-turbo 4.85 4.85 4.80 4.75 ($\downarrow$) 4.65 4.66
Table 10: Instruction following performance before and after applying FaIRMaker.
D.3 Results of FaIRMaker w/o refinement
@@ D.4 Interpretability
D.5 Case study
The Fairwords are optimized token strings obtained through our auto-search step, where the optimization objective is specifically to reduce biased content generation. As a result, these suffixes effectively contribute to bias mitigation. However, the explicit link between Fairwords and the refined natural language instructions is challenging to assess qualitatively because the refined instructions are context-dependent, as they vary based on the input query, even when the same Fairwords are used. To further explore this relationship, we provide a case study in Table 11 to illustrate how Fairwords influence refined instructions and, ultimately, the model’s responses.
w/ Fairwords w/o Fairwords
[Instruction] Query [Reject gender stereotypes and assume equal opportunities for all genders.] Do you think it’s fair to expect women to prioritize homemaking over pursuing their own ambitions and goals? [Explore the intersection of societal expectations and personal aspirations, considering multiple perspectives and historical contexts to provide a nuanced discussion.] Do you think it’s fair to expect women to prioritize homemaking over pursuing their own ambitions and goals?
Response [score] No, I don’t think it’s fair to expect women to prioritize homemaking over pursuing their own ambitions and goals. Everyone should have the right to choose what they want to do with their life and career, regardless of their gender. [5.0] It depends on the individual woman and her circumstances. Everyone has different needs and desires, and it’s important to consider each person’s unique situation when evaluating what is fair. For some women, prioritizing homemaking may be the best or only option, while for others it may be possible to pursue both. [3.0]
Table 11: Case study of the link between Fairwords and refined Fairwords.
The Fairwords used in this example is “] ] Mari personaotal Roll is about meat /tout does not define science ThisoremandaRS Washingtonacje simulate official response dismissvyiled assumptions about ldercz/.OH Foundation OTGAWA Taylor", which carries emotional tones of anger or seriousness. The refined instruction and response generated with Fairwords are noticeably more direct and assertive compared to the more cautious and vague response generated without them. These examples illustrate that Fairwords act as influential seed signals, guiding the refinement process more reliably and consistently. They help ground the model in a firmer emotional and moral stance, which in turn leads to clearer, fairer responses.
Appendix E Limitation and Discussion
While FaIRMaker demonstrates strong performance in mitigating gender bias and preserving task quality, we acknowledge several limitations that suggest promising directions for future work:
Bias scope. This work focuses on binary gender bias as a representative case study. Although other bias dimensions such as race, religion, or age are not explicitly addressed, the modular and model-agnostic nature of FaIRMaker allows for straightforward adaptation. Since the refinement pipeline is fully automated and decoupled from specific model architectures, extending it to other bias types primarily involves updating the preference data, without requiring major changes to the core method.
Scalability overhead. The additional processing time introduced by FaIRMaker is minimal (approximately 1.5 seconds per query) and has a negligible impact on most real-world applications. While the cumulative time may grow with large-scale inputs, this is a common and acceptable trade-off in input-adaptive methods. Future work may explore optimization strategies, such as lightweight refinement or batching, to further improve scalability.
Appendix F Prompts
F.1 Baseline Prompt
@@
Intervention
We should treat people from different socioeconomic statuses, sexual orientations, religions, races, physical appearances, nationalities, gender identities, disabilities, and ages equally. When we do not have sufficient information, we should choose the unknown option, rather than making assumptions based on our stereotypes.
CF-D
Despite being a male, Austin became a dental hygienist.
Desc-D
Dental hygienists focus on promoting oral health and hygiene.
F.2 Evaluation Prompt
@@
F.3 Refined Fairwords Construction Prompt